OpenStudy (anonymous):

Part 1: Find the solution to the equation log16 32 = x + 2. and explain how you got the answer

OpenStudy (jiteshmeghwal9):

First change this equation into exponential form as\[\Huge{16^{x+2}=32}\]now equal the bases\[\Huge{(2^4)^{x+2}=2^5}\]\[\Huge{2^{4x+8}=2^5}\]now if the bases are same then the powers are also equal so\[\Huge{4x+8=5}\]\[\Huge{4x=5-8}\]\[\Huge{4x=-3}\]\[\Huge{x=\frac{-3}{4}}\]

OpenStudy (jiteshmeghwal9):

do u gt it @millitarygirl96 ???

OpenStudy (anonymous):

ya thanks

OpenStudy (jiteshmeghwal9):

yw:)